# Devising PoPStat: A metric bridging population pyramids with global disease mortality

### v1

@@ Abstract
Understanding the relationship between population dynamics and disease-specific mortality is central to evidence-based health policy. This study introduces two novel metrics, PoPDivergence and PoPStat, one to quantify the difference between population pyramids and the other to assess the strength and nature of their association with the mortality of a given disease. PoPDivergence, based on Kullback-Leibler divergence, measures deviations between a country’s population pyramid and a reference pyramid. PoPStat is the correlation between these deviations and the log form of disease-specific mortality rates. The reference population is selected by a brute-force optimization that maximizes this correlation. Utilizing mortality data from the Global Burden of Disease 2021 and population statistics from the United Nations, we applied these metrics to 371 diseases across 204 countries. Results reveal that PoPStat outperforms traditional indicators such as median age, GDP per capita, and Human Development Index in explaining the mortality of most diseases. Noncommunicable diseases (NCDs) like neurological disorders and cancers, communicable diseases (CDs) like neglected tropical diseases, and maternal and neonatal diseases were tightly bound to the underlying demographic attributes whereas NCDs like diabetes, CDs like respiratory infections and injuries including self-harm and interpersonal violence were weakly associated with population pyramid shapes. Notably, except for diabetes, the NCD mortality burden was shared by constrictive population pyramids, while mortality of communicable diseases, maternal and neonatal causes and injuries were largely borne by expansive pyramids. Therefore, PoPStat provides insights into demographic determinants of health and empirical support for models on epidemiological transition. By condensing the multi-dimensional population pyramid to a scalar variable, PoPDivergence allows us to examine the relationship of any variable of interest with the underlying population structure. These metrics have the potential to advance future global health and population research.
1Department of Electrical and Electronic Engineering, Faculty of Engineering, University of Peradeniya.
2Department of Community Medicine,3Department of Pharmacology,
Faculty of Medicine, University of Peradeniya.
@@ 1 Introduction
Understanding the social and economic determinants of mortality causes is crucial for developing effective healthcare policies, interventions, and prevention programs. To this end, many studies have explored factors that help explain the variation in mortality causes. For instance, the Human Development Index (HDI), a composite metric combining life expectancy, education, and income, is used to examine cancer mortality patterns across regions [1]. These factors are often challenging to measure directly due to underlying, unobserved circumstances, commonly referred to as latent variables. Latent variables have been known to impact multiple causes of death. Among them, demographic variables are particularly significant, as they can provide valuable insights into observed mortality trends.
The effects of demographic and cultural variations, as well as how different groups responded to various diseases, have been examined using factors such as median age, income level and population density, all of which are valuable for policymaking[2, 3, 4]. Furthermore, demographic parameters have been utilized to track disease progression within communities, and correlating these parameters with disease spread has become a common practice for both communicable[5, 6, 7, 8] and non-communicable[9, 10, 11] diseases. A recent study by Carnegie et al indicates that high population density appears to be associated with high mortality rates of various non-communicable diseases (NCD) in western developed countries[12]. Additionally, research by Li et al. demonstrates that epidemic sizes are strongly linked to population density[13], while a study by Pardhan et al. shows a negative relationship between GDP per capita and the increase in new COVID-19 cases across European countries. These scalar variables, particularly the median age, are derived from a country’s population distribution. However, due to the median being an average-like statistic, it does not provide insights into the spread or shape of the distribution, which is essential for understanding the population structures. Consequently, the characteristics associated with different population pyramid shapes, expansive and constrictive forms, are not captured by the median age, which fails to reflect demographic variation. Regardless, the vast usage of median age proves the applicability of the information from the population distribution even in its most basic form is warranted[14, 15, 16].
The population pyramid breaks down the population according to gender: male, female, and age group. As shown in figure 1, the most prominent population pyramid structures are expansive, stationary, and constrictive. Expansive structures, also referred to as progressive, are considered youthful populations with a broad base indicating high birth rates. Stationary structures represent populations that are stable in size, with birth and death rates approximately balancing each other, leading to a roughly uniform distribution across age groups. Constructive structures, also referred to as regressive, are considered aged populations with a narrower base indicating low birth rates. Nevertheless, for brevity, these quantitative data of each country will be referred to as the population pyramid[17].
@@
The evolution of the pyramid shape is a result of historical events[18] such as major policy changes, natural phenomena, and political events as these factors affect the birth and death rate of countries. For example, the Middle East population doubled from 25 million to 54 million between 2005 and 2015 as the number of migrant individuals and families increased due to economic opportunities and China’s pyramid shrank inwards for the age cohorts around 50, showing the slowing of population growth during the Great Chinese famine. After the famine, the population started increasing, but about 30 years ago, the pyramid started thinner again with the “one-child policy”[19, 20] enacted in 1979. Hence, the population pyramid reflects national policies[18] and how said policies affect the country’s living conditions.
Apart from the derivative variables of the population pyramid, such as median age, a correlation exists between the economic variables and the pyramid shape[21]. For example, a macroeconomist can analyze the pyramid shape of a country to explore the relationship between population growth and GDP per capita growth. Similarly, a health economist might examine how local population composition impacts the demand for long-term care facilities. Besides, countries with different population compositions (but similar in other respects) will allocate their resources differently. For example, an inverted pyramid implies the government could negotiate more health service payments. Consequently, their GDPs will also differ, as countries with stronger government control over healthcare costs may allocate resources differently, leading to varying economic outcomes. Understanding these connections is crucial for policymakers seeking to address the specific needs of a population based on its demographic structure.
These economic and societal correlations highlight the population pyramid’s potential as an informative representation of demographic structures. However, there is a need for novel quantitative approaches that can extend beyond traditional measures like median age or population growth rates to capture the intricate relationships between demographic compositions and their impacts on diverse domains. This gap necessitates the development of new scalar variables that are capable of representing demographic variability in a more robust and actionable manner.
As mentioned earlier, the literature highlights the instrumental role of the population pyramid in policy-making [22, 23, 24]. Policymakers could utilize the proposed variable or mechanism to devise more informed policies. Moreover, the proposed mechanism can be adapted to identify countries or regions within a country, provided sufficient data is available, that share similar socioeconomic standards and enable benchmarking against their disease severity or vulnerability. The ability to correlate demographic structures with health outcomes not only advances our understanding of population dynamics but also lays the groundwork for comparative studies with other demographic and development indicators.
To build on this foundation, we consider related research that investigates similar connections between population metrics and health outcomes. A similar approach was taken by Zhu et al. as they studied the relationship between coronary heart disease and the Human Development Index (HDI)[25]. Additionally, as stated in [13], population density is also considered in investigating the correlation with severity measures. Similarly, numerous socioeconomic indicators exist; however, based on the literature, median age, population density, GDP per capita, HDI, and life expectancy were chosen for comparison in this study due to their strong association with population demographics and their ability to explain variations in health outcomes[26, 27, 28, 29, 30, 31, 32, 33]. Later in the study, the effectiveness of the proposed scalar variables was assessed by comparing their correlation results with these widely used socioeconomic indicators.
Figure 2: Difference in population distribution for similar (a) Median Age, (b) GDP per Capita, (c) Life Expectancy, (d) Population Density, and (e) Human Development Index
Figure 3: Difference in (a) Median Age, (b) GDP per Capita, (c) Life Expectancy, (d) Population Density, and (e) Human Development Index for similar population distributions
In Figure 2, a few examples illustrate the significance of the population pyramid from the other commonly used variables: median age, population density, GDP per capita, HDI, and life expectancy for statistical analysis. In Figure 3, We have provided the population pyramids of the countries with similar indices for the other variables but with disparate demographic distributions and vice versa. The difference suggests that the alternative variables are affected by factors other than the population demographics.
Population pyramids that are similar in shape more often than not correspond to similar median ages. Nonetheless, the converse of that statement does not hold water, as countries with similar median ages might have dramatically different population pyramids, as shown in Figure 2(a). This is because the median age is indifferent to any symmetrical changes in the population pyramid. Accordingly, it is conclusive that countries with closely matching population pyramids have similar demographic profiles, while such a claim is not necessarily valid when using the median age as a parameter.
This study explores the relationship between population demographics and disease severity, aiming to understand how variations in demographic distributions influence national mortality rates for different diseases. By examining these correlations, we seek to uncover patterns that can provide deeper insights into the impact of population structures on mortality causes. Bearing this goal in mind, we introduce two novel scalar variables: PoPDivergence and PoPStat. PoPDivergence is designed to encourage the use of a more informative representation of demographic distributions, while PoPStat quantifies the strength of the correlation between PoPDivergence and a severity measure for a given disease. The methodology used to generate these variables will be properly discussed in the following sections. This study examines the application of these two scalar variables as explanatory variables for various mortality causes, facilitating the correlation of demographic distribution variances with national mortality rates for different diseases. The variable was applied to all-cause mortality data for 2021, specifically focusing on recorded deaths per million. This novel approach aims to statistically quantify the age distribution, or the shape of the population pyramid, in relation to mortality data for various cases of death.
The results highlight how these metrics capture latent information in population pyramids, offering insights into the demographic, social, and cultural factors that influence mortality. The performance of PoPStat was shown to be superior to other demographic and socio-economic indicators, especially for non-communicable diseases (NCDs), and it performed better than traditional development metrics like HDI and GDI in explaining disease mortality patterns. It also provides new perspectives on the Epidemiological Transition Model (ETM)[34] by showing how NCDs align with constrictive population pyramids, while communicable diseases are more associated with expansive pyramids. The findings underscore the utility of PoPDivergence and PoPStat as tools for understanding the complex interplay between population structure and health outcomes, offering potential for future research in other areas of social, behavioral, and economic studies.
Further, the association of the population pyramid to the economy allows allocating funds[35] optimally to critical sectors identified as vulnerable zones based on their demographics now linked more effectively to different causes of death. This signifies the importance of the PoPStat as it can correlate to various socioeconomic and healthcare attributes through the structure of the population pyramid because these socioeconomic and healthcare standards directly impact the death rate. While medical research is about the prevention and eradication of diseases, analysis with an essential tool such as a population pyramid is effective in understanding the factors that affect different rates for mortality causes other than medical.
@@ 2 Methods
2.1 Mortality Data
Mortality data was retrieved from the Global Burden of Disease (GBD) Study 2021 conducted by the Institute for Health Metrics and Evaluation (IHME) which includes estimates for 371 diseases and injuries across 204 countries. The causes were systematically organized within a four-tiered hierarchical classification progressively increasing in specificity which are mutually exclusive and collectively exhaustive at each level. Level 1 consisted of three broad aggregate categories (communicable, maternal, neonatal, and nutritional (CMNN) diseases; non-communicable diseases; and injuries). Level 2 disaggregated these categories into 22 cause clusters, while Level 3 further expanded to 175 causes, comprising 132 specific causes and 43 clusters derived from Level 4 causes. For this analysis, causes that had fatal outcomes were examined up to Level 3.
2.2 Population Data
Population data was obtained from the Population Division, Department of Economic and Social Affairs, United Nations as of World Population Prospects 2024. Population pyramids for each country were then constructed for 2021 by organizing the population data by five-year age groups. The data for each age category were then normalized by dividing the population in each group by the total population, effectively representing the proportion of the population for each gender across five-year age intervals.
@@ 2.3 Constructing the PoPDivergence
Population pyramids are multidimensional representations of population structure. To establish a scalar parameter that can be correlated with demographic metrics such as mortality rates, we propose PoPDivergence, a distance measure between population pyramids. This approach involves calculating the divergence from a pre-selected reference population pyramid, quantifying the extent to which one population pyramid can be inferred from the reference, and evaluating the deformation between the structural shapes of two population distributions.
PoPDivergence is a Kullback-Leibler (KL) divergence-based metric to measure the relative difference between two probability distributions. Assuming two population distributions $P$ and $Q$, with $P_{(i)}$ and $Q_{(i)}$ representing the proportion of the population in the $i^{th}$ age category, and $Q$ serving as the reference population, we calculate the PoPDivergence as mentioned in equation 1,
@@
which quantifies the relative entropy between the two distributions. Specifically, the metric measures the information-theoretic distance from the reference population P to the comparative population Q.
2.4 PoPDivergence Interpretation
Figures 4 and 5 illustrate the quantification of differences between population pyramids using two distinct reference populations; the Central African Republic and Japan. A PoPDivergence of zero indicates no Kullback-Leibler (KL) divergence between population pyramids, which occurs when comparing the reference country with itself. For all other countries, the PoPDivergence reflects the KL divergence between their respective population pyramids and the reference population. When using the Central African Republic as the reference population—a country in the early stages of demographic transition—an increase in PoPDivergence reveals the progression from a progressive to a regressive population pyramid. Conversely, with Japan as the reference population—representing a country in the later stages of demographic transition—an increasing PoPDivergence indicates a shift from a regressive to a progressive population structure. The interpretation of PoPDivergence changes must always be contextualized by the specific reference population.
0
(a) Central African Rep.
0.009
(b) Chad
0.186
(c) India
0.643
(d) Poland
0.664
(e) Malta
Figure 4: PoPDivergence for the reference of Central African Republic
0
(a) Japan
0.026
(b) Finland
0.079
(c) Russia
0.320
(d) Bangladesh
0.518
(e) Kenya
Figure 5: PoPDivergence for the reference of Japan
We consider PoPDivergence as the explanatory variable for disease-specific mortality rates (number of deaths due to a given cause per one million of the population of a given year), to compute the correlation. Since PoPDivergence’s distance computation is not symmetric, we establish the reference population using the objective function below in 2. This maximizes the correlation between cause-specific crude mortality rate in the natural log scale and PoPDivergence.
@@ 2.4 PoPDivergence Interpretation
where $\Omega$ is the set of all countries included in the study, $S$ is the cause-specific crude death rate, and PoPDivergence($\omega$) is the distance between two population pyramids with $\omega$ as the reference country. We used brute force searching to establish the reference population pyramid because there is no closed-form solution to the above optimization problem. It is computationally economical given the discrete nature of the variables involved. This method is what we introduce as reference country tuning, which identifies the population demographic most strongly correlated with a specific disease’s death rates.
@@ 2.5 Constructing the PoPStat
PoPStat quantifies the relationship between cause-specific death rates and the structure of population pyramids across countries. To derive this metric, we first determine a reference country for a given cause through reference country tuning. Subsequently, for each country, the PoPDivergence from the reference country is calculated and stored in vector X. The corresponding cause-specific death rates, transformed to the natural logarithm scale, are stored in vector Y. PoPStat is then defined as the Pearson correlation coefficient between X and Y, as calculated using Equation 3, thereby representing the relationship between PoPDivergence (reference) and cause-specific mortality.
@@
where, $x_{i}$ and $y_{i}$ are the PoPDivergence value and natural log of the number of deaths per million for the $i^{th}$ country, respectively, and $x,y$ are the mean values of vectors $X$ and $Y$. In this study, we will consistently present Popstat in the format (PopStat value, reference country, p-value).
@@ 2.6 Comparison of PoPStat with other indicators
When comparing the PoPStat metric with other potential indicators, mentioned in the introduction also, we began by identifying the most commonly used demographic indices in the literature, specifically life expectancy, median age, and population density for the countries analyzed. Furthermore, to assess the extent to which population pyramids capture latent information about social and economic determinants of disease mortality, we included Gross Domestic Product (GDP) per capita as a macroeconomic indicator and the Human Development Index (HDI) as a well-established composite development measure for the relevant countries. Next, we calculated the Pearson correlation coefficients between each of these indices and cause-specific death rates to compare their associations of mortality against PoPStat.

### v2

@@ Abstract
Background Traditional demographic indicators offer a limited view of a country’s population structure and may not capture demographic influences on mortality patterns comprehensively. We aimed to bridge this gap by developing novel scalar metrics that condense the information in population pyramids and assess their association with disease specific mortality.
Methods Country specific population pyramids were constructed using the United Nations World Population Prospects 2024, while mortality data for 371 diseases across 180 countries were extracted from the Global Burden of Disease Study 2021. We then developed two metrics: PoPDivergence, which quantifies the difference between a country’s population pyramid and an optimized reference pyramid using Kullback Leibler divergence, and PoPStat, which is the correlation between PoPDivergence and cause specific mortality rates.
Findings Non communicable diseases (NCDs) showed a strong PoPStat of –0.84 (optimized reference: Japan, p<0.001) with mortality concentrated to constrictive pyramids. Communicable, maternal, neonatal, and nutritional diseases had a moderate PoPStat of 0.50 (Singapore, p<0.001) with mortality linked to expansive pyramids. Injuries exhibited a weak PoPStat (0.291, Singapore, < 0.001). In more granular analyses, NCDs like neurological disorders and neoplasms, communicable diseases (CDs) like neglected tropical diseases, and other infections showed a strong PoPStat. Nevertheless, NCDs like diabetes, cirrhosis, and other chronic liver diseases, CDs like respiratory infections and tuberculosis carried a weak PoPStat indicating minimal influence of population pyramid on their mortality.
Interpretations This study, its devised metrics, demonstrates the degree to which the mortality of different diseases is bound to the underlying population structure and reveals what type of population pyramids will carry the highest mortality attributed to those diseases.
Funding This study received no external funding.
Keywords Population Pyramids, Demographic Transition, Epidemiology, PopDivergence, PoPStat
Codespace The full source code and analytic scripts for this study are available in the GitHub Codespace repository at https://github.com/Buddhi19/DevisingPoPStat.git.
Research in context
Evidence before this study
To assess the current state of evidence, we systematically searched PubMed, Scopus, and Google Scholar on December 5, 2024, using different combinations of the keywords, “mortality”, “population pyramid”, “median age”, “life expectancy”, "demographic" and “correlation” with no restrictions placed on language or publication date. Previous research has studied associations between well known demographic indicators, such as median age, population density, and life expectancy, and various causes of mortality. In relation to population pyramids, we identified only narrative syntheses that assessed their impact on cause specific mortality. Most quantitative work on population pyramids have focused on clustering patterns and population projections. There remains a significant research gap in developing scalar metrics that can robustly quantify the influence of the entire population pyramid shapes on mortality outcomes.
Added value of this study
This study was a novel attempt to quantitatively analyze the association between the shape of population pyramids and the type of mortality causes. We do so by constructing population pyramids from the most recently available population datasets and utilizing mortality data from a database with comprehensive coverage. We introduced two new metrics, PoPDivergence and PoPStat, that broaden our understanding of demographic determinants of disease specific mortality. They also offer empirical support to the demographic and epidemiological transition models.
Implications of all the available evidence
By summarizing information rich population pyramids into a scalar metric PoPDivergence opens a new avenue in research to assess association of any social, behavioural or economic variable with population pyramid morphology. The ability of PoPDivergence to locate stage of each country in the demographic transition allows it to capture population change, that happens across generations, within a single time point. This facilitates research on demographic transition without requiring longitudinal data over decades. Meanwhile, PoPStat provides a framework in epidemiology to study how underlying demographic structure impacts mortality of different diseases. As implications for practice, at the national level, PoPStat enables reliable predictions on disease specific mortality using population pyramid shapes, in contexts where mortality data may be missing or insufficient for a specific disease. Our findings also inform health policymakers of the types of diseases they need to prioritize given the stage of demographic transition they are in, which can guide resource allocation and planning in health. For global health, these metrics can introduce a novel at risk approach among countries based on the vulnerabilities of their population pyramid shapes for specific diseases.
@@ 1 Introduction
Globally, approximately 60 million deaths occur per year, with their causes often being tied to socioeconomic disparities[1]. Understanding social and economic determinants of mortality is therefore crucial for developing effective healthcare policies and preventive interventions. Consequently, many studies have explored factors influencing mortality variations, with socioeconomic indicators like income level, Socio Demographic Index (SDI), and Human Development Index (HDI) being central to understanding mortality patterns[2, 3]. Additional social determinants, such as access to healthcare, employment conditions and living standards, have also been studied extensively for their impact on mortality outcomes[4, 5].
Demographic characteristics of the population, median age, life expectancy, and population density have been frequently utilized to study cause specific mortality[6, 7]. Research has shown how populational attributes influence the spread and severity of both communicable and non-communicable diseases[8, 9]. High population density has been linked to increased transmission rates of infectious diseases[10], while aging populations exhibit higher mortality rates for conditions such as cardiovascular diseases and neurodegenerative disorders[11]. However, conventional demographic indicators have inherent limitations. Using metrics related to the central tendency of population distribution is insufficient for understanding nuanced demographic influences on mortality patterns. Figures 1 illustrates these discrepancies between demographic indicators and their respective population distributions underscoring the limitations of traditional metrics.
(a) Difference in median age for similar population distribution.
(b) Difference in population distribution for similar median age.
(c) Difference in life expectancy for similar population distribution.
(d) Difference in population distribution for similar life expectancy.
(e) Difference in population density for similar population distribution.
(f) Difference in population distribution for similar population density.
Figure 1: Differences in median age, life expectancy, and population density, together with the corresponding variations in population structure.
A more comprehensive approach to understanding demographic structure is through population pyramids which represents the distribution of a population by age and sex. In population pyramids, age is typically divided into five year strata that are stacked vertically in an ascending order. Each strata represents the numbers of both males and females in that age group[12]. As shown in Figure 2, population pyramids are categorized into expansive (progressive), stationary, and constrictive (regressive) forms, reflecting youthful, stable, and aging populations, respectively[13]. This unique visualization of population pyramids allows them to capture the entire age-sex spectrum of a country overcoming the limitations of traditional demographic variables. Nevertheless, no methodological approach has yet been proposed in literature to condense these information rich structures into comparable scalar variables.
@@
The evolution of the population pyramid from an expansive to a constrictive shape is best explained by the landmark Demographic Transition Model (DTM) introduced by Warren Thompson. Though this transformation is directly driven by birth, death, and migration rates as proposed by the DTM, it is also closely associated with policy changes, economic shifts, natural disasters and, more importantly, epidemiological trends[14]. Building on the DTM, Abdel Omran’s Epidemiological Transition Model (ETM) posited a direct relationship between transformations in population structures and transitions in cause-specific mortality [15]. Therefore, among demographic variables, examining the association of population pyramids with cause specific mortality becomes a priority. However, a thorough literature review revealed that this relationship has been mainly assessed through narrative syntheses[16] suggesting population pyramids have been underutilized in quantitative epidemiology, limiting their role in mortality studies.
Therefore, this study has two primary objectives, (1) to develop a quantitative method for extracting and condensing age-sex distribution data from population pyramids into a scalar variable, and (2) to employ this variable to investigate the association between disease-specific mortality rates and underlying demographics.
@@ 2 Methods
2.1 Data Sources and Preprocessing
Mortality data for 371 diseases and injuries across 180 countries were obtained from the Global Burden of Disease Study 2021 (IHME). Population data were sourced from the United Nations World Population Prospects 2024, with 2021 country specific population pyramids constructed using five year age groups and normalized to reflect gender specific proportions. Causes were organized in a four tiered hierarchy, and for our analysis, fatal outcomes up to Level 3 were considered.
@@ 2.2 Constructing the PoPDivergence
Population pyramids are multidimensional representations of population structure. However, a scalar metric is needed to correlate them with mortality rates. We propose PoPDivergence as shown in 1, a Kullback Leibler (KL) divergence based measure to quantify the structural deviation of a population pyramid from an optimized reference population.
@@
Here, $P$ and $Q$ denote population distributions, with $P_{(i)}$ and $Q_{(i)}$ representing the proportion of the population in the $i^{th}$ age category, and $Q$ serving as the optimized reference population.
This measures the information theoretic distance between demographic structures. Figures 3 illustrates PoPDivergence using the Central African Republic (CAR) and Japan as reference populations, deliberately chosen to represent the extremes of the demographic transition. A zero value indicates identical distributions. Contextual interpretation depends on the reference population pyramid.
PopDivergence 0 0.009 0.186 0.643 0.664
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Central African Rep. Chad India Poland Malta
PopDivergence 0 0.026 0.079 0.320 0.518
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Japan Finland Russia Bangladesh Kenya
Figure 3: Comparison of PoPDivergence values across countries. Top row: Central African Republic as reference, showing transition from progressive to regressive population pyramids. Bottom row: Japan as reference, indicating movement toward progressive structures.
2.3 Reference Country Tuning
Since, the PoPDivergence values depend on a pre selected reference country, to maximize correlation with cause specific mortality rates (expressed as the natural log of deaths per million), we optimize the reference selection using the objective function 2,
@@ 2.3 Reference Country Tuning
where $\Omega$ is the set of countries, and $S$ is the mortality rates. Given the absence of a closed form solution, brute force search identifies the reference pyramid with the strongest correlation to mortality. This method is what we introduce as reference country tuning, which enables the identification of the optimized population demographic(out of 180 countries), which is most strongly associated for a given disease’s mortality rate.
@@ 2.4 Constructing the PoPStat
PoPStat quantifies the association between cause specific mortality and the structure of population pyramids. For each cause, a reference country is first selected through reference tuning. Then, the PoPDivergence from this reference is calculated for every country (stored in vector X), while the natural logarithms of the corresponding cause specific death rates are stored in vector Y. PoPStat is then defined as the Pearson correlation coefficient between X and Y.
@@
Here, $x_{i}$ denotes the PoPDivergence of the $i^{\text{th}}$ country, and $y_{i}$ is the natural logarithm of its cause specific mortality rate. The terms $\bar{x}$ and $\bar{y}$ represent the means of vectors $X$ and $Y$, respectively.
In this study, we will consistently present Popstat in the format (PopStat value, reference country, p-value).
@@ 2.5 Comparison of PoPStat with other indicators
To benchmark PoPStat, its association with cause specific mortality was compared with traditional demographic indices, life expectancy, median age, and population density, as well as socioeconomic measures such as GDP per capita and the Human Development Index (HDI). Pearson correlation coefficients between each indicator and the cause specific death rates were calculated to evaluate their relative predictive strengths.